OpenStudy (anonymous):

e^2x -7e^2x+12=0 there are 2 answers they want both

OpenStudy (barrycarter):

Replacing e^(2x) with y, we have: y - 7*y + 12 = 0 or y=2 Thus e^(2x) = 2 We can rewrite e^(2x) as (e^x)^2, so we know that: e^x = +- Sqrt[2] = +-Sqrt[2^(1/2)] Taking the log, the answers are: x = Log[2]/2 and x= Log[2]/2 + I*Pi (the second answer is complex)

OpenStudy (anonymous):

webwork says that isnt the right answer

OpenStudy (barrycarter):

wolframalpha.com seems to agree with me: http://fwd4.me/0m5L